The material for the bolt shown in the angled joint above has an ultimate tensile strength of 500 MPa and a shear strength of 300 MPa. The diameter of the bolt is 8 mm. Determine the factor of safety in operation.


1
Expert's answer
2021-01-20T17:46:05-0500


Assume that the bolt is perpendicular to the inclined plane. Then the force can be resolved into components that are normal and tangential to the plane. The two components are

"F_n=F\sin\theta, F_t=F\cos\theta."

The area of the 8 mm bolt is "A=\dfrac{\pi d^2}{4}=\dfrac{\pi (8\times10^{-3})^2}{4}=16\pi\times10^{-6}m^2"

The normal and tangential stresses are 


"\sigma_n=F_n/A=\dfrac{F\sin\theta}{16\pi}MPa"

"\sigma_t=F_t/A=\dfrac{F\cos\theta}{16\pi}MPa"

The factors of safety are ("F" in "N" )


"f_n=\dfrac{\sigma_n^{allow}}{\sigma_n}=\dfrac{500MPa}{\dfrac{F\sin\theta}{16\pi}MPa}=\dfrac{8000\pi}{F\sin\theta}"

"f_t=\dfrac{\sigma_t^{allow}}{\sigma_t}=\dfrac{300MPa}{\dfrac{F\cos\theta}{16\pi}MPa}=\dfrac{4800\pi}{F\cos\theta}"

Let "F=8kN=8\times10^3N, \theta=50\degree." Then the factors of safety are


"f_n=\dfrac{8000\pi}{8000\sin50\degree}=4.1"

"f_n=\dfrac{4800\pi}{8000\cos50\degree}=2.9"